I - AMFE A uniform beam of length 2 m and a mass of 45 kg is mo 15° angle form horizontal, supported on one end by a hinge connecting it to a wall, and on the other end by a rope that is at a 15° angle to horizontal as shown. What is the tension rope?
Answer #1


In equilibrium net torque = 0

m*g*L/2*cos15 - T*cos15*L*sin15 - T*sin15*L*cos15 = 0

m*g*L/2*cos15 - T*L*sin30 = 0


T*sin30 = m*g*cos15/2


T*sin30 = 45*9.8*cos15/2

Tension T = 426 N